OpenStudy (anonymous):

1/3 (x-2) = x + 4 HELPPP

OpenStudy (ankit042):

where is the problem? you have to solve for x..cross multiply and continue

OpenStudy (anonymous):

i have to multiply x-2 by 1/3 right?

OpenStudy (anonymous):

@ankit042

OpenStudy (texaschic101):

distribute the 1/3 through the parenthesis

OpenStudy (ankit042):

or you can multiply both the side with 3

OpenStudy (texaschic101):

true

OpenStudy (nurali):

\[\frac{ 1 }{ 3(x-2) }=x+4\] or \[\frac{ 1 }{ 3 }(x-2)=x+4\] which one?

OpenStudy (nurali):

reply @melissalovee!

OpenStudy (anonymous):

the second one @Nurali

OpenStudy (nurali):

ok

OpenStudy (ankit042):

@Nurali please let her do it on her own...you can guide her to reach the answer

OpenStudy (nurali):

\[\frac{ 1 }{ 3 }(x-2)=x+4\] \[(x-2)=3(x+4)\] \[x-2=3x+12\] \[3x-x=-12-2\] \[2x=-14\] \[x=\frac{ -14 }{ 2 }\] \[x=-7\]

OpenStudy (anonymous):

@Nurali thanks a lot. Is that a rule that instead of distributing the 1/3 you can just multiply the other side by 3?

OpenStudy (nurali):

multiply both sides by 3.

OpenStudy (anonymous):

Oh so you are getting rid of the 1/3 basically

OpenStudy (nurali):

Yes

OpenStudy (ankit042):

You can do by either way. This is basic algebra, You might find this useful http://www.dragonboxapp.com/

OpenStudy (anonymous):

Thanks @ankit042. Learned this back in 7th grade so its hard for me to remember it all.

OpenStudy (ankit042):

np :)

OpenStudy (anonymous):

yeah that is right